# In-Context Learning Dynamics with Random Binary Sequences

### v1

@@ 5.3 Distinguishing Formal Languages from Randomness
GPT-3.5 sharply transitions between behavioral patterns, from generating pseudo-random values to generating non-random sequences that perfectly match the formal language (Fig. 8). We observe a consistent pattern of formal language learning in GPT-3.5 generating random sequences where predictions $p(y|x)$ of depth $d\geq 4$ are initially random with small $|x|$, and have low $p(y\in C|x)$ where $C$ is a given concept. This follows whether the prompt describes the process as samples from “a weighted coin” or “a non-random-algorithm”. We also find sharp phase changes in GPT-3.5 behavioral patterns in Randomness Judgment tasks across 9 binary concepts (Fig. 4). These follow a stable pattern of being highly confident in that the sequence is Random (high $p(y=\mathtt{random}|x$) when $x$ is low, up to some threshold of context at which point it rapidly transitions to being highly confident in the process being non-random. Transition points vary between concepts, but the pattern is similar across concepts (also see Fig. 1314 in Appendix).
6 Conclusion
@@
We find strong patterns of random number Generation and Judgment in LLMs, including an emergent ability to generate low-variance pseudo-random binary sequences that can be partly explained with a memory-constrained moving window model, and partly as memorization and compression. We also find striking behavioral phase changes when GPT-3.5 learns to distinguish simple formal language concepts from randomness. This supports theories of ICL as Bayesian model selection and as under-specified program induction, and demonstrates how LLM behavioral patterns can change non-linearly with ICL as a function of context data. Future Cognitive Interpretability work may characterize concept learning dynamics underlying other LLM capabilities, including those with applications in usability and safety. If ICL is indeed analogous to hypothesis search in human cognition (or, equivalently, under-specified program induction) where new concepts can be learned in-context or computed in text generation trajectories, with no changes to model weights, a compelling research direction would be to further understand the features and limits of in-context learning dynamics using theories and tools from computational cognitive science.
@@ Appendix A Related Work
Language model evaluations  Our work resembles evaluation benchmarks such as BIG-Bench (Srivastava et al., 2022) that use behavior alone to evaluate LM understanding and reasoning. However, as described later, the domain of subjective randomness is fundamentally different in that there is no “correct” answer. Linguistic probing attempts to characterize the structure of LM representations, but unlike our work, is a function of hidden unit activations rather than output behavior.
@@ Appendix B Additional Experimental Details
System
Your responses will only consist of comma-separated "Heads" and "Tails" samples.
Do not repeat the userś messages in your responses.
User
[…]
Assistant
@@
Appendix C Formal Language Learning, by Depth
@@
Figure 11: Predictive distribution $p(y|x)$ trees with $d=6$ for concept $C=(011)^{n}$ with $|x|\in\{6,12,18\}$. Since $|x|$ is increasing by the same depth as the tree $\Delta_{|x|}=d=6$, the transition from generating pseudo-random numbers to deterministically repeating $011$ is visibly apparent. Also see Figure 8.
@@
Appendix D Randomness Judgments
Figure 13: Randomness judgments across GPT models for 9 concepts Also see Fig. 4.
@@ Appendix D Randomness Judgments
(Fig 13) text-davinci-003 shows a stable pattern of being highly confident (high token probability) in the process being random up to some amount of context $|x|$, at which point it rapidly transitions to being highly confident in the process being non-random, with transition points varying substantially between concepts. chat-gpt-3.5-instruct does not go through a stable high-confidence random period like text-davinci-003, and stable high-to-low confidence dynamics are observed for only a subset of concepts. The majority of earlier GPT models (text-davinci-002, text-davinci-001, text-curie-001, text-babbage-001) show no ’formal language learning’, at all. However, surprisingly OpenAI’s smallest available GPT model text-ada-001 shows S-shaped in-context learning dynamics, with the peak close to .5 instead of 1.0 as in text-davinci-003. Additionally, the learning dynamics and transition points for all concepts appear nearly identical, approximately at $|x|=50$, and some concepts show less stable "non-random" patterns for larger $|x|$.
@@
Figure 14: Randomness Judgment $p(y=\mathtt{random}|x)$ dynamics for each concept tested, for text-davinci-003 and gpt-3.5-turbo-instruct
Appendix E Random Sequence Generation by GPT Model

### v3

@@ 5.3 Distinguishing Formal Languages from Randomness
Figure 8: GPT repeats certain sub-sequences more often than chance. We find that GPT-3.5 repeats certain sub-sequences more often than is predicted by our Window Average model. While the Window Average model generates fewer unique sub-sequences than a Bernoulli process, this does not account for the bias in GPT-3.5 (text-davinci-003) to repeat many of the same sub-sequences. This disparity increases with longer sub-sequences (Fig. 26).
@@
GPT-3.5 sharply transitions between behavioral patterns, from generating subjectively random values to generating non-random sequences that perfectly match the formal language (Fig. 5). We observe a consistent pattern of formal language learning in gpt-3.5-turbo-instruct-0914 generating random sequences where predictions $p(y|x)$ of depth $d\geq 4$ are initially random with small $|x|$, and have low $p(y\in C|x)$ where $C$ is a given concept. This follows whether the prompt describes the process as samples from “a weighted coin” or “a non-random-algorithm”. ICL curves for longer concepts are smoother for text-davinci-003 and do not converge to $p(y|x)\approx 1$ as they do for gpt-3.5-turbo-instruct-0914. We do not explore the reason for this distinction or when this breakdown occurs; however, in App. G we analyze longer concepts in Randomness Judgment experiments, which are more cost-efficient.
We also find sharp phase changes in GPT-3.5 behavioral patterns in Randomness Judgment tasks across 9 binary concepts (Fig. 4). These follow a stable pattern of being highly confident in that the sequence is Random (high $p(y=\mathtt{random}|x$) when $x$ is low, up to some threshold of context at which point it rapidly transitions to being highly confident in the process being non-random. Transition points vary between concepts, but the pattern is similar across concepts (also see Fig. 1718 in Appendix). These transitions support our theory of ICL as model selection, or discrete hypothesis search, rather than model averaging, as in Fig. 2.
6 Discussion
@@
We find strong patterns of random number Generation and Judgment in LLMs, including an emergent ability to generate low-variance, ‘subjectively’ random binary sequences, that can be partly explained with a memory-constrained moving window model. This behavior can also partly be explained as simple memorization and compression, but we observe algorithmic behavior that seems to go beyond mere parroting. We also find striking behavioral phase changes when GPT-3.5 learns to distinguish simple formal language concepts from randomness. This supports our framework of understanding ICL as Bayesian model selection and as under-specified program induction, and demonstrates how LLM behavioral patterns can change non-linearly with ICL as a function of context data. Our work is an initial step in a direction which may prove to be highly impactful: if new capabilities can suddenly, rather than gradually, emerge after a single additional data point is presented in-context, this presents major challenges to AI safety and LLM interpretability. LLM capabilities may suddenly emerge or disappear with new prompts, or with more ‘shots’, and LLM behavior may suddenly and dramatically shift over the course of a single user interaction. These phase changes may be common in ICL, unobserved by research in few-shot learning that averages over many samples when considering in-context learning curves, instead of considering individual learning trajectories as we do. Future work may characterize concept learning dynamics underlying other LLM capabilities, including applications in usability and safety (App. C). If ICL is analogous to hypothesis search in human cognition, a compelling research direction would be to further understand the defining features and limits of in-context learning dynamics using theories and tools from computational cognitive science.
Reproducibility Statement
All code and data used for this paper are available at https://github.com/ebigelow/ICL-Random-Binary. In Appendix L we replicated our results across a selection of open-source LLMs available through the HuggingFace Transformers API (Llama 2, Mixtral, Tulu 2). We found similar results to GPT-3.5+ on Randomness Generation tasks varying $p(Tails)$ with LLMs with 50+ billion parameters (Fig. 32 &  33) and similar results on Randomness Judgment with tulu-2-dpo-70b (Fig. 30). As of January 2024, the primary LLM analyzed in this paper text-davinci-003 is no longer publicly accessible.
Acknowledgements
ESL’s time at University of Michigan was partially supported via NSF under award CNS-2008151.
@@ Appendix A Related Work
Language model evaluations  Our work resembles evaluation benchmarks such as BIG-Bench (Srivastava et al., 2022) that use behavior alone to evaluate LM understanding and reasoning. However, as described in the main text, the domain of subjective randomness is fundamentally different in that there is no “correct” answer. Linguistic probing attempts to characterize the structure of LM representations, but unlike our work, is a function of hidden unit activations rather than output behavior.
@@ Appendix B Additional Experimental Details
System Your responses will only consist of comma-separated "Heads" and "Tails" samples.
Do not repeat the user’s messages in your responses.
@@
[…]
Figure 9: Consistent tokenization with ‘Heads, Tails, …’ compared to ‘010 …’ With comma-separated lists of ‘Heads’ and ‘Tails’ flips, tokens consist of ‘Heads’, ‘T’, ‘ails’, ‘,’, as well as occasional erroneous tokens such as ‘Head’. With comma-less ‘0’ and ‘1’ tokens, the GPT tokenizer chunks many binary sub-sequences into individual tokens.
We chose the format “Heads”, “Tails”, … specifically to avoid the alternate interpretation you describe, of some patterns merely being tokenization artifacts (Fig. 9), since flips with shorter string formats, such as 0s and 1s with no comma separating them, often get chunked into different tokens (“01”, “$0$”, etc.). We specify ‘1000 random samples’ in our Generation prompt to avoid issues with GPT generating too few tokens, where for example if the ‘50 samples’ is specified, less than 50 flips will be generated in some cases.
[…]
Appendix C Cognitive Interpretability
Figure 10: Cognitive Interpretability in the context of prior work Cognitive Interpretability studies in-context learning dynamics in LLMs, positing theories of the latent concepts enabling behavioral capabilities. It is a middle ground between behavioral benchmarks, which treat models as black boxes and evaluate hit-or-miss accuracy, and mechanistic interpretability, which studies toy models and training loss dynamics, analogous to how cognitive science is a middle ground between behaviorist psychology and neurobiology.
Cognitive Interpretability (Fig. 10) is a middle ground between shallow test-set evaluation benchmarks on one hand (Srivastava et al., 2022; Saparov & He, 2022; Min et al., 2022; Lanham et al., 2023; Ganguli et al., 2023; Bowman et al., 2022; Kadavath et al., 2022; Perez et al., 2022; Prystawski & Goodman, 2023; Turpin et al., 2023) and mechanistic neuron- and circuit-level understanding of pre-trained model capabilities on the other (Nanda et al., 2023; Goh et al., 2021; Geva et al., 2020; Belinkov, 2022; Li et al., 2022; Wang et al., 2022; Chughtai et al., 2023; Gurnee et al., 2023; Foote et al., 2023; Lubana et al., 2023; Lieberum et al., 2023; Barak et al., 2022; Liu et al., 2022b; Zhong et al., 2023) (also see App. C). Unlike related work of feature interpretability in vision neural networks, we analyze ICL and text generation dynamics over different input data $x$ and outputs $y$, and unlike mechanistic work studying ICL, we observe learning dynamics in black box input-output behavior of state-of-the-art LLMs.
This is inspired by cognitive science work where learning dynamics are observed in humans, and these learning dynamics are explained with computational theories and probabilistic programming models. For example, children 3.5–5 years old learning to count undergo a dramatic conceptual shift from knowing the meanings of only a few number words (“one”, “two”) to a full inductive understanding of counting, which can be modeled as Bayesian model selection with a simplicity prior over models (Piantadosi et al., 2012). Another example is children learning intuitive theories of physics and social reasoning, where theories may spontaneously shift as ‘a new concept is acquired’. This can be explained, similar to our models, as a shift in the posterior space, which influences hypothesis search (Ullman et al., 2012). The work of Ullman et al. (2012) is also of note, since they make a similar point to the neural network Quantization hypothesis Michaud et al. (2023) – that sharp discontinuities (‘phase changes’) are common in learning, but averaging over samples and individuals turns these discontinuities into a smooth learning curve – but in the context of children’s intuitive theory learning and hypothesis search over probabilistic programs. Similar theories are also proposed about theory change in adults, for example, in scientific concepts dramatically shifting as paradigms change (Carey, 2000; Nersessian, 2010).
In-Context Learning and emergent capabilities in LLMs present unique and novel challenges and opportunities for AI interpretability, as well as for cognitive science. Computational cognitive science provides rich theory and myriad domains which can be used for evaluating LLM learning and behavioral dynamics. Our approach is distinct from approaches such as Binz & Schulz (2023) which evaluate LLMs on cognitive psychology benchmarks, since such prior work examines single-token probabilities but does not consider, as we do, in-context learning dynamics or applications of computational theories of cognition (Tenenbaum, 1998; Tenenbaum et al., 2011; Ullman & Tenenbaum, 2020; Piantadosi, 2021) to understanding LLM capabilities. Our approach is also distinct from using LLMs as models of human cognition. Though we use theories and models from cognitive science to understand LLMs, we do not assume that what the LLM does should be human-like.
While we propose a new term, Cognitive Interpretability, we opt to provide a loose definition rather than a strict one. The purpose of this term is to emphasize a missing piece of the LLM interpretability puzzle, which is not given as much attention as evaluation benchmarks or mechanistic circuit-level understanding. Mechanistic interpretability is sometimes described as the neuroscience of deep learning. We suggest a role for the cognitive science of deep learning, distinct from creating new evaluation benchmarks, which seeks to understand the high-level algorithms underlying complex patterns of behavior and learning.
Appendix D Bayesian Model Selection Example
This section further explains the illustrative example presented in Fig. 2.
[…]
Figure 11: (Left) Normalized hypothesis posteriors from Fig. 2 Left. These are normalized so that probability of all hypotheses sums to one $\sum_{h}p(h|x)=1$, i.e. $p(h=\text{Bernoulli}|x)+p(h=\text{Non-Random}|x)=1$. (Right) The soft model selection curve shown in Fig. 2 Right, with varying temperature parameter $\tau$ interpolating $p_{\tau}(y|x)$ between Model Averaging ($\tau=1$, yellow line) and Model Selection ($\tau\rightarrow\infty$, dark blue line).
In the right side of Figure 2, we compare Model Averaging $p(y|x)=\sum p(y|h)\ p(h|x)$ to Model Selection $h^{*}=\operatorname*{argmax}_{h}p(h|x),\ \ p(y|x)=p(y|h^{*})$. With model selection, at the point when the the two hypothesis posteriors cross, the predictive distribution $p(y|x)$ suddenly shifts as the non-random hypothesis becomes selected for $h^{*}$ instead of the random one. This example illustrates how model selection, or maximum likelihood hypothesis search, can explain sharp phase changes in model behavior according to the predictive distribution $p(y|x)$.
Since the phase change in-context learning curves we observe are smooth, as well as those observed during neural network training and mechanism formation, we also show a ‘soft’ model selection curve. The soft model selection curve interpolates between model averaging and model selection (see Fig. 11, Right):
$p_{\tau}(y|x)=\sum_{h}p(y|x)\big{(}\frac{p(h|x)^{\tau}}{\sum_{h^{\prime}}p(h^{\prime}|x)^{\tau}}\big{)}$
where $\tau$ is a temperature parameter such that when $\tau=1$, the function $p_{\tau}(y|x)$ is equal to the model averaging definition above, and when $\tau$ grows arbitrarily large, $p_{\tau}(y|x)$ approaches the model selection definition above.
Temperature $\tau$ in this model is distinct from temperature in LLMs, which leads models to greedy decode as $\tau\rightarrow 0$, i.e. $p(y|x)=\max_{h}p(y|h)\ p(h|x)$. This maximizes over the posterior predictive distribution, which includes both $p(y|h)$ and $p(h|x)$, whereas in our example, $h^{*}$ only maximizes over the hypothesis posterior $p(h|x)$.
Appendix E Formal concept learning is not merely local bias
In our formal concept learning experiments, we found similar results whether the prompt specified generating samples ‘that may be from a fair coin with no correlation, or from some non-random algorithm’ or only ‘from a fair coin, with 50% probability of Heads and 50% probability of Tails’. This raises the question of whether the LLM is simply forgetting the prompt, and has a local text bias that leads it to be distracted by a long repeating pattern, instead of following the original task specification, as in neural text degeneration (Holtzman et al., 2019).
To test this alternative hypothesis, we used a real-world task used to evaluate few- and zero-shot learning in LLMs: the Massive Multitask Language Understanding (MMLU) dataset (Hendrycks et al., 2020). Each MMLU task includes a single multiple choice question, along with 4 multiple choice answers: A, B, C, and D. MMLU tasks are organized into subjects, intended to cover a broad range of common academic and professional subjects. We tested with two questions randomly selected from each of 5 subjects in MMLU: College Physics, Elementary Mathematics, High School Biology, High School Psychology, Philosophy.
Our experimental setup was such that we followed the MMLU original prompt format, which ends with the text ‘Answer: ____’. Although MMLU provides options for few-shot learning, we used zero context shots. We appended a sequence of flips ‘Heads, Tails, Heads, Tails, …’ from the concept $(\texttt{01})^{n}$ to the end of this prompt, similar to how our experiments use a prompt format: Q: …  A: Heads, Tails, ….
[…]
As shown in Fig. 12, GPT-3.5 (text-davinci-003) shows dramatically different learning patterns in this MMLU setup, compared with our Randomness Generation results. In most MMLU questions, GPT is extremely robust against local text bias, maintaining a high probability of staying on-task and generating MMLU tokens even after hundreds of tokens (note: each flip, comma, and space together constitute 2-3 tokens) following a simple repeating pattern are appended to the prompt.
Figure 12: GPT-3.5 demonstrates a task bias, returning to answer MMLU questions after many distractor flips In these experiments, we take MMLU questions and append sequences of ‘distractor flips’, repeating samples from $(\texttt{01})^{n}$, before querying the LLM for next-token probability. Individual lines are shown for flip tokens (‘Heads’, ‘Tails’, ‘T’, ‘ails’) compared with MMLU answer tokens (‘A’, ‘B’, ‘C’, ‘D’). Even with many distractor tokens, GPT-3.5 converges back to answering the MMLU question, demonstrating that our formal concept learning results for the Generation task are not merely a local text bias where the LLM forgets the original prompt.
The primary goal of this experiment was not to explore ICL dynamics in this MMLU distractor experiment, however, we do note striking patterns for different MMLU tasks. Though a few questions seem to demonstrate stable phase changes, sharply converging to high-confidence of generating flip tokens, we note that all but one (College Physics 1) of these converge to a state where MMLU tokens still have 5-20% probability. This probability means that, when temperature=1.0 and stochastic decoding is used to sample completions, nearly all completions eventually converge back to the MMLU task, ending with an answer instead of more flips.
[…]
Although not shown here, we also found similar results appending distractor flips to prompts from GSM8k (text-davinci-002 and -003), a simple mathematical benchmark commonly used to evaluate chain-of-thought reasoning. GPT’s task bias appeared stronger in GSM8k, but was harder to evaluate, since GSM8k does not consist of multiple choice questions like MMLU.
@@ Appendix E Formal concept learning is not merely local bias
Appendix F Formal Language Learning with Varying Prediction Depth
In our main text, for formal language learning with Generation tasks, we estimate the predictive probability $p(y\in C\ |\ x)$ by enumerating all possible binary sequences $y$ up to some depth $d$ and querying GPT for next-token prediction probabilities. In this appendix, we show results of this analysis with varying prediction depth $d$.
@@ Appendix F Formal Language Learning with Varying Prediction Depth
A notable takeaway of Figures  13,  14 is that with shallower depths ($d<4$), we do not observe as stable ICL dynamics. Our predictive probability method is similar to a simple Markov Chain, where a greater depth $d$ similar to a higher Markov Chain order $k$, enables more expressive representation of $p(y|x)$. If $d=2$, for example, then we only consider the 4 conditional probabilities $p(y=0|x=0)$, $p(y=0|x=1)$, $p(y=1|x=0)$, $p(y=1|x=1)$. With our methods of computing the predictive distribution, there are only three strings $y^{*}$ such that $y^{*}\in C$ for any depth, where at a depth $d$ there will be $2^{d}$ possible strings $y$. We also note that the predictive distributions $p(y|x)$ are increasingly close together, suggesting that with $d>6$, the trends we observe will only change slightly.
Figure 15: Predictive distribution $p(y|x)$ trees with $d=6$ for concept $C=(011)^{n}$ with $|x|\in\{6,12,18\}$. Since $|x|$ is increasing by the same depth as the tree $\Delta_{|x|}=d=6$, the transition from generating subjectively random numbers to deterministically repeating $011$ is visibly apparent. Also see Figure 5.
@@
In Figures  15,  16, we show predictive distribution $p(y|x)$ trees, as in Figure 5 in the main text. Figure  15 should be read left-to-right, and Figure  16 should be read top-to-bottom. For example, in  15, the middle figure begins $x=\texttt{011011011011}$ where the left figure ends ($x=\texttt{011011011011},y=\texttt{011011011011}$), conditioned on the outcome sequence $y=\texttt{011011}$. The right-most figure similarly begins where the left one ends.
Like Borges’ Garden of Forking Paths  (Bottou & Schölkopf, 2023), at each step in context (i.e. at some $|x|$ for a given $x$), LLM text generation presents the user with branching paths of possible token sequences. Some of these paths may lead to very different outcomes, in our case, either deterministically repeating some pattern or generating random flips. The domain of binary sequences lets us deeply explore and easily visualize a minimal version of this.
Appendix G Randomness Judgments
Figure 17: Randomness judgments across OpenAI GPT models for 9 concepts  Also see Fig. 4. OpenAI models ordered from smallest/least fine tuned to largest/most: text-ada-001 (??), text-babbage-001 (1B params + SFT), text-curie-001 (6.7B + SFT), text-davinci-001 (175B + SFT), text-davinci-002 (175B + SFT), text-davinci-003 (175B + RLHF-PPO), gpt-3.5-turbo-instruct (??). These OpenAI model details were previously publicly available and have been archived at: https://archive.is/IXxGm. Note: technically text-davinci-002 is a ‘GPT-3.5’ model, but in our text we use ‘GPT-3.5+’ to refer to text-davinci-003 and later GPTs.
@@ Appendix G Randomness Judgments
(Fig 17) text-davinci-003 shows a stable pattern of being highly confident (high token probability) in the process being random up to some amount of context $|x|$, at which point it rapidly transitions to being highly confident in the process being non-random, with transition points varying substantially between concepts. chat-gpt-3.5-instruct does not go through a stable high-confidence random period like text-davinci-003, and stable high-to-low confidence dynamics are observed for only a subset of concepts. The majority of earlier GPT models (text-davinci-002, text-davinci-001, text-curie-001, text-babbage-001) show no ’formal language learning’, at all. However, surprisingly OpenAI’s smallest available GPT model text-ada-001 shows S-shaped in-context learning dynamics, with the peak close to .5 instead of 1.0 as in text-davinci-003. Additionally, the learning dynamics and transition points for all concepts appear nearly identical, approximately at $|x|=50$, and some concepts show less stable “non-random” patterns for larger $|x|$.
@@
Figure 18: Randomness Judgment $p(y=\mathtt{random}|x)$ dynamics for additional concepts tested, for text-davinci-003 (Left) and gpt-3.5-turbo-instruct (Right).
Longer Formal Concepts
Testing text-davinci-003 with longer concepts (Fig.19), many concepts still follow a pattern of sharply transitioning from high confidence in $p(y=\mathtt{random}|x)$ to high confidence in $p(y=\mathtt{non}\text{-}\mathtt{random}|x)$. With increasingly long concepts, this pattern further degrades. This degradation is not surprising, for two reasons: (a) large concepts are more costly to represent, and so may either be more difficult for the LLM to learn, or the learning curves that show these patterns may be much wider (e.g. $|x|$ ranging from 0 to 1000 instead of 0 to 100, as we use). (b) Large concepts are less frequent in the training data, and so may be more prone to memorizing samples. When only a few samples are available, neural networks should be expected to memorize data points and context (Feldman, 2020), since there is not enough context to represent that pattern as a generic formal concept.
Figure 19: Randomness Judgment with longer formal concepts Similar to Fig. 4, we test Randomness Judgment with GPT-3.5 using longer repeating concepts. From left to right, concept lengths are: 8, 10, 12. Legends list colors corresponding to individual concepts.
Random Sequence Baseline
[…]
While there is a higher overall average $p(Random|x)$ (blue line) for small $|x|$, with larger $|x|$ the average $p(Random|x)$ remains flat around $0.6$. Contrast this with the formal concepts in Fig.4, where with large $|x|$, $p(Random|x)$ drops to nearly $0.0$
[…]
Appendix H Random Sequence Generation by GPT Model
In Figures  21,  21, we show flip sequences $y$ from the Randomness Generation task, with various $p(\text{Tails})$ super-imposed on the same plot, with different colors according to the specified $p(\text{Tails})$. In Fig. 23, we compare the mean sequence probability $1/N\sum_{i}\overline{y^{(i)}}$, averaged over all GPT output sequences $y^{(i)}$, to the prompt-specified probabilities $p(\text{Tails})$.